We are given three coins. One has heads on both faces, the second has tails on both faces, and the third coin has a head on one face and a tail on the other face. We choose one coin at random, toss it, and observe that the result is heads. What is the probability that the opposite face is tails?
Answer #1

This is a case of Conditional Probability.

Let there be two events A and B. Event A denotes the probability of getting a head from the toss of three given coins. Event B is the event of the other side of the resultant coin in first toss being tails.

P(A|B)=P(AintersectionB)/P(B)

We know that P(AintersectionB)=1/3 (Since there’s only one coin that can both give heads and be tails on the opposite side).

And, P(B)=2/3 (Since there are 2 coins that can show tails on the opposite face)

Therefore, P(A|B)=(1/3)/(2/3)=1/2

.. Or, there is a very easy way to look at this answer. Given that you already got heads, now your sample space is only reduced to 2 coins (since the one with both sides tails can’t give heads). We are asked for the probability of the coin being the one with opposite sides. Therefore,

P=(No.of favourable Outcomes)/(Total no. of outcomes)=1/2